OpenStudy (anonymous):

Given: f(x)= 3x^2-3x and g(x) = x^2 +x -12 find the domain f/g

OpenStudy (anonymous):

set the denominator = 0 and solve. your answer will be all real numbers except...

OpenStudy (anonymous):

I got x=4 and x=-3, but 4 is ok so only 3 or -3

OpenStudy (anonymous):

ok I did it right ayay!!!

OpenStudy (anonymous):

\[x^2+x-12=0\] \[(x+4)(x-3)=0\] \[x=-4\] or \[x=3\] all real numbers except -4 and 3

OpenStudy (anonymous):

think you factored correctly but solved incorrectly yes?

OpenStudy (anonymous):

why not -4< i know it factors that way bu tplug it in?

OpenStudy (anonymous):

your factors are (x+4)(x-3) so your solutions are -4 and 3 yes?

OpenStudy (anonymous):

yes so would the answer be (-infinity,-4) u (-4,3) U (3, infinity)

OpenStudy (anonymous):

\[(x+4)(x-3)=0\] \[x+4=0\] \[x=-4\] or \[x-3=0\] \[x=3\]

OpenStudy (anonymous):

yes you got it!

OpenStudy (anonymous):

which is a rather wordy way of saying all numbers except -4 and 3

OpenStudy (anonymous):

yipee! the only part I dont get is why -4, I know is factors down to that, but it doesnt make the equations numerator =0?